Use Laplace transformation to solve differential equation.
*+ 4y = e, y(0) = druge dt - dºg(0)
Answer #1

given differential equation dy +4832 0(0)-decod 9600 i. yall +44 = et, y(o)= y CO )= y (0)=0 applying Laplace transform on bcomparing 82 coefficients 0 = A + B = 3 = -A B = -2 companing s coefficients 0 - - 13+ =) c= 13- S-1 Ict) u- - [3Ct) 3 = 16 ]